The following data for an imaginary economy.

MPC = b = 0.8 , autonomous consumption = $50 M , Investment = $100 M , Government spending = $80 M, Transfer payments = $40 M, Taxes = $12 M

Using this data calculate:

a. The level of national income

b. The level of Saving
Answer #1

a. At equilibrium, Y = C + I + G

Y = 50 + 0.8(Y+40-12) + 100 + 80

Y = 252.4 + 0.8Y

Y = 1262

National income = 1262

b. C= 50 + 0.8(1262+40-12) = 1082

National savings = Y - C - G

= 1262 - 1082 - 80 = 100